Grade 12Electrostatics

A small ball of mass 2 × 10-3 kg having a charge of 1 μC is suspended by a string of length 0.8 m. Another identical ball having the same charge is kept at the point of suspension. Determine the minimum horizontal velocity which should be imparted to the lower ball so that it can make complete revolution in a vertical circle. Take g = 10 m/s2

Shakti Nirmal
8 Years agoGrade 12

1 Answer

Arun
8 Years ago

 m .g - k . q/ l= m. v2 / l            .... (1)

             m.v02/2 + k.q2/l = 1/2.m.v2 + k.q2/l + m.g.2l ...........(2)

             From 1 & 2, v0 = 5.86 m/s